# spacing=nonfrench

### v1

@@ 6 Appendix
Let $\Gamma$ be a Jordan curve with $\infty\notin\Gamma$, and $\Omega_{i,e}$ the connected components of $\overline{\mathbb{C}}\setminus\Gamma$ as above. In all this work we have treated $\Omega_{i}$ and $\Omega_{e}$ similarly: in doing so we have overlooked the fact that $\Omega_{e}$ is unbounded, thus requiring an extra argument. The purpose of this appendix is to fill this gap.
Without loss of generality, suppose that $\Gamma$ is included in the unit disk $\mathbb{D}$ with 0 being an interior point. The map $z\mapsto 1/z$ transforms $\Omega_{e}$ into a bounded domain $\Omega^{\prime}_{i}$ with boundary $\Gamma^{\prime}=1/\Gamma$; $\Omega_{i}^{\prime}$ contains 0 which corresponds to $\infty$ in $\Omega_{e}$. We say $u$ is harmonic in the domain $\Omega_{e}$ with $\infty$ being an interior point if $u(1/z)$ is harmonic in the bounded domain $\Omega_{i}^{\prime}$. Let $p>1$ and $s\in(0,1)$ as usual and consider $u\in\mathcal{B}_{p,p}^{s}(\Omega_{e})$. Let $v:\Omega_{i}^{\prime}\to\mathbb{C}$ being defined by $v(z)=u(1/z)$. By a change of the variable $z=1/w$ we conclude by the Koebe distortion theorem (2.1) that
@@
where the implicit constant is a universal one. We have the following
@@
Let $r<1/4$ and $R_{1},R_{2}$ be such that $\frac{1}{2}+r<R_{1}<R_{2}<1-r$. For any $\zeta$ such that $|\zeta|<r$ the annulus $A_{\zeta}=\{R_{1}<|z-\zeta|<R_{2}\}$ is included in the annulus $\{\frac{1}{2}<|z|<1\}$. By the mean value property for harmonic functions we have, for $|\zeta|<r$,
$|\nabla v(\zeta)|=\frac{1}{\pi R_{i}^{2}}\iint_{|z-\zeta|<R_{i}}\nabla v(z)dxdy,\quad i=1,2$
@@
Noting (6.1) we can conclude by using Hölder inequality that $|\nabla v|$ is uniformly bounded in the disk $|\zeta|<r$ by a constant depending only on $\Gamma$ multiplied by the norm of $u$ in $\mathcal{B}_{p,p}^{s}(\Omega_{e})$. This estimate implies that
@@
In order to control the rest of the integral giving the norm of $v$:
$\iint_{\Omega^{\prime}_{i}\setminus(|z|<r)}|\nabla v(z)|^{p}d(z,\Gamma^{\prime})^{(1-s)p-1}dxdy,$ (6.3)
@@
Theorem 6.2.
If $\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$, and the identity operator from $\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)$ onto $B_{p,p}^{s}(\Gamma)$ is a bounded isomorphism, then the same assertions hold for the unbounded component $\Omega_{e}$ of $\overline{\mathbb{C}}\setminus\Gamma$, and in particular, the transmission operator from $\mathcal{B}_{p,p}^{s}(\Omega_{i})$ onto $\mathcal{B}_{p,p}^{s}(\Omega_{e})$,
$\mathcal{B}_{p,p}^{s}(\Omega_{i})\to\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)\to\mathcal{B}_{p,p}^{s}(\Omega_{e}),$
is a bounded isomorphism.
Proof.
Let us use the above proposition to prove that $\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$. If $u\in\mathcal{B}_{p,p}^{s}(\Omega_{e})$ then $v\in\mathcal{B}_{p,p}^{s}(\Omega^{\prime}_{i})$ so that $v\in\mathcal{B}_{p,p}^{s}(\Omega_{i}^{\prime}\!\to\!\Gamma^{\prime})=B_{p,p}^{s}(\Gamma^{\prime})$ from which it follows easily that $u\in B_{p,p}^{s}(\Gamma)$. ∎
Statements and Declarations. The authors declare that there is no conflict of interest, and there is no data associated to this work.
Acknowledgments. This work is supported by the National Natural Science Foundation of China (Grant No. 12271218 and 12571083) and the University of Orléans. The authors also warmly thank the Poincaré Institute in Paris for its hospitality through the framework of the program "Research in Paris". Special thanks also to Gérard Bourdaud for his help and for interesting discussions.

### v2

@@ 6 Appendix
Let $\Gamma$ be a Jordan curve with $\infty\notin\Gamma$, and $\Omega_{i,e}$ the connected components of $\overline{\mathbb{C}}\setminus\Gamma$ as above. In all this work, we have treated $\Omega_{i}$ and $\Omega_{e}$ similarly: in doing so we have overlooked the fact that $\Omega_{e}$ is unbounded, thus requiring an extra argument. The purpose of this appendix is to fill this gap.
Without loss of generality, suppose that $\Gamma$ is included in the unit disk $\mathbb{D}$ with 0 being an interior point. The map $z\mapsto 1/z$ transforms $\Omega_{e}$ into a bounded domain $\Omega^{\prime}_{i}$ with boundary $\Gamma^{\prime}=1/\Gamma$; $\Omega_{i}^{\prime}$ contains 0 that corresponds to $\infty$ in $\Omega_{e}$. We say $u$ is harmonic in the domain $\Omega_{e}$ with $\infty$ being an interior point if $u(1/z)$ is harmonic in the bounded domain $\Omega_{i}^{\prime}$. Let $p>1$ and $s\in(0,1)$ be as usual and consider $u\in\mathcal{B}_{p,p}^{s}(\Omega_{e})$. Let $v:\Omega_{i}^{\prime}\to\mathbb{C}$ be defined by $v(z)=u(1/z)$. By a change of variables $z=1/w$ we conclude by the Koebe distortion theorem (2.1) that
@@
where the implicit constant is universal. We have the following
@@
Let $r<1/4$ and $R_{1},R_{2}$ be such that $\frac{1}{2}+r<R_{1}<R_{2}<1-r$. For any $\zeta$ such that $|\zeta|<r$ the annulus $A_{\zeta}=\{R_{1}<|z-\zeta|<R_{2}\}$ is included in the annulus $\{\frac{1}{2}<|z|<1\}$. By the mean value property for harmonic functions, we have, for $|\zeta|<r$,
$\nabla v(\zeta)=\frac{1}{\pi R_{i}^{2}}\iint_{|z-\zeta|<R_{i}}\nabla v(z)dxdy,\quad i=1,2$
@@
Noting (6.1) we can conclude using Hölder inequality that $|\nabla v|$ is uniformly bounded in the disk $|\zeta|<r$ by a constant depending only on $\Gamma$ multiplied by the norm of $u$ in $\mathcal{B}_{p,p}^{s}(\Omega_{e})$. This estimate implies that
@@
In order to control the rest of the integral, giving the norm of $v$:
$\iint_{\Omega^{\prime}_{i}\setminus\{|z|\leq r\}}|\nabla v(z)|^{p}d(z,\Gamma^{\prime})^{(1-s)p-1}dxdy,$ (6.3)
@@
Statements and Declarations. The authors declare that there are no conflicts of interest and that there are no data associated with this work.
Acknowledgments. This work is supported by the National Natural Science Foundation of China (Grant Nos. 12271218 and 12571083) and the University of Orléans. The authors also warmly thank the Poincaré Institute in Paris for its hospitality through the framework of the program "Research in Paris". Special thanks also to Gérard Bourdaud for his help and for interesting discussions.